OpenStudy (anonymous):

(x^2 - 2x)/(2x - 4) Reduce to lowest terms, and please explain!

OpenStudy (campbell_st):

factorise the numeratorand denominator (x^2 - 2x) = x(x - 2) -------- ------- (2x - 4) = 2(x - 2) cancelling the common factor will leave the solution

OpenStudy (anonymous):

x/2?

OpenStudy (campbell_st):

yes... nice and simply

OpenStudy (anonymous):

Thank you so much!